Kabbed:

A basketball player gets 2 free-throw shots when she is fouled by a player on the opposing team. She misses the first shot 40% of the time. When she misses the first shot, she misses the second shot 5% of the time. What is the probability of missing both free-throw shots?

Kabbed:

I'll use P(A) • P(B) = P(A ∩ B)

Kabbed:

@Vocaloid

Kabbed:

@Hero I need help I don't know what to do next I will use P(A) • P(B) = P(A ∩ B) to start but what do I do next

Kabbed:

@lowkey @Narad

Kabbed:

P(A) • P(B) = P(A ∩ B)

Kabbed:

its blank

lowkey:

oh

lowkey:

The answer to your question is 40/100 * 5/100 0.4 * 0.05 = 0.02 = 2%

Kabbed:

I need to use P(A) • P(B) = P(A ∩ B)

Kabbed:

help me use it, it should be easy because now we know the answer

lowkey:

Oh it's just p(0.4) times p(0.05)=P(0.4*0.05)

Kabbed:

what does ∩ mean

lowkey:

You got sass.

Kabbed:

p(0.4) times p(0.05)=P(0.4*0.05)

lowkey:

ya

Kabbed:

p(0.4) times p(0.05)=P(0.4*0.05) p(0.2)

Kabbed:

is that whats next

Kabbed:

p(0.4) times p(0.05)=P(0.4*0.05) and then p(0.2) or do we do another step before we get 0.2

lowkey:

thats ok

Kabbed:

thank you so much